# ode]python

### v2

@@ Abstract
Recent advances in machine learning (ML) have accelerated progress in calibrating and operating quantum dot (QD) devices. However, most ML approaches rely on access to large, representative datasets designed to capture the full spectrum of data quality encountered in practice, with both high- and low-quality data for training, benchmarking, and validation, with labels capturing key features of the device state. Collating such datasets experimentally is challenging due to limited data availability, slow measurement bandwidths, and the labor-intensive nature of labeling. QDFlow is an open-source physics simulator for multi-QD arrays that generates realistic synthetic data with ground-truth labels. QDFlow combines a self-consistent Thomas-Fermi solver, a dynamic capacitance model, and flexible noise modules to simulate charge stability diagrams and ray-based data closely resembling experiments. With an extensive set of parameters that can be varied and customizable noise models, QDFlow supports the creation of large, diverse datasets for ML development, benchmarking, and quantum device research.
@@ 1 Introduction
Among the various quantum computing platforms, quantum dots (QDs) stand out for their scalability potential, compact size, and long coherence times [Burkard21-SSQ]. Operating QD devices, however, remains a formidable challenge, with complexity growing rapidly—often exponentially—as the number of qubits increases. Recent advances in integrating machine learning (ML) with quantum device operation have begun to mitigate these difficulties, offering promising automated control and calibration strategies. For example, ML algorithms have been developed for the fabrication [1, 2], characterization [3, 4], tuning [5, 6, 7, 8, 9, 10, 11, 12], and gate virtualization [13, 14] of QD devices.
Developing robust ML models requires access to large and diverse datasets representative of the multi-dimensional parameter space typical of QD devices. Crucially, for supervised ML applications, these datasets must also include metadata that identifies key features, such as the global state (i.e., the number of QDs formed), the charge configuration, and the type of transition lines present. Unfortunately, large volumes of high-quality experimental data can be challenging to obtain as companies and research groups often keep such data proprietary [15]. Limited measurement bandwidth in real-world experiments also constrains the efficient exploration of the entire high-dimensional parameter space in a reasonable time. Generating accurate feature labels for publicly available data is a labor-intensive and time-consuming task that can produce subjective and potentially erroneous labels. Physics-based simulations offer a practical solution: they enable the generation of arbitrarily large datasets while providing direct access to the ground-truth charge states, thereby simplifying the labeling process needed for ML training.
@@
There are several open-source QD device simulators [16, 17, 18], which rely on the constant capacitance model, treating the array of QDs and their associated electrostatic gates as nodes in a network of fixed capacitors. Additionally, Ref. [18] allows the capacitances to vary with respect to the number of charges $n$ by introducing an $n$-dependent correction to the capacitance matrix. In contrast, in QDFlow the capacitance parameters are physics-informed, obtained directly from the self-consistent Thomas-Fermi solution rather than imposed heuristically. All key physical observables—such as current, charge states, and sensor readouts—are derived from a capacitance model constructed based on the computed charge density $n(x)$. This ensures that the capacitances evolve dynamically with gate voltages, yielding a more realistic description of device behavior. Furthermore, QDFlow allows for modeling regions with low barriers between dots, leading the dots on either side to combine into a single centralized dot. Finally, QDFlow incorporates a flexible noise module, enabling the addition of experimentally relevant effects such as thermal broadening, charge offset drift, and voltage fluctuations. These features make the simulated data qualitatively comparable to experimental measurements while maintaining full access to the ground truth labels required for ML applications. Building on QFlow—a legacy implementation of the QD simulator that applied the Thomas-Fermi approximation to model charge densities and stability diagrams [19, 20]QDFlow extends these methods into a flexible, open-source framework that integrates physics-informed capacitance modeling with realistic noise processes tailored for ML applications.
@@
QDFlow has already demonstrated its utility in advancing ML-driven QD research. The legacy version of the simulator was used to generate the QFlow-lite dataset [21, 20], which enabled the training of several ML models for global state recognition [19, 8, 11]. The dataset also supported the development of a novel classification framework for simple high-dimensional geometrical structures, known as the ray-based classification (RBC) framework [22]. The expanded dataset, QFlow 2.0: Quantum dot data for machine learning [21], generated using the Thomas-Fermi solver with integrated realistic noise processes, further advanced ML-based approaches to QD tuning. In particular, models trained with data from the QFlow 2.0 dataset have been successfully applied to tasks such as data quality assessment[23], physics-informed RBC and ray-based navigation in 1D QD arrays [12], and the development of a full virtualization stack for 1D and 2D QD arrays [13]. These early successes highlight the value of physics-informed synthetic datasets for accelerating the development of automated control tools for QD systems.
@@
QDFlow is publicly available for download from the Python Package Index, with the source code released under the GNU General Public License and comprehensive documentation and tutorials to support new users included in the QDFlow GitHub repository [24]. The library includes full type hints for all classes and functions to ensure clarity, maintainability, and ease of extension.
@@ 2 Physics simulation
QDFlow has three main modules: simulation, noise, and generate, and one utility module, distribution, as depicted in Fig. 1. The simulation and noise modules are part of the physics package. The core physics-based engine of the simulator is contained in the simulation module. It uses a Thomas-Fermi solver to find the stable charge configuration and sensor output of a particular QD device defined by a set of physical parameters. The QDFlow Thomas-Fermi solver was first introduced in Ref. 20, but has since been refined and extended within QDFlow to improve flexibility, physical relevance, and integration with downstream ML workflows. The PhysicsParameters dataclass, which is used to initialize the simulation, specifies over twenty parameters governing the properties of the QD device. These parameters include both material characteristics and device-specific features such as gate geometry and positioning. Importantly, the gate voltages—experimentally relevant control knobs—are explicitly included among the simulation inputs. By sweeping these voltages, QDFlow produces the final outputs: 2D CSDs and 1D rays, directly mirroring the tuning procedures used in real QD experiments.
@@
To generate data with QDFlow, the user first chooses whether to run the default configuration or adjust the distributions and ranges over which physics parameters are randomized. Next, they create one or more sets of randomized device parameters, and for each device, generate a CSD using the functions in the generate module. Once the physics parameters have been specified, an instance of the ThomasFermi class is instantiated. This class serves two main purposes: first, it solves for the charge density function $n(x)$; and second, it uses $n(x)$ to construct a capacitance model and compute physical quantities such as the device’s charge state and sensor response. QDFlow then runs the physics simulation for every pixel in each diagram and compiles the results into a CSDOutput dataclass, which is returned to the user. By repeating this process over a range of gate voltages, QDFlow generates data that can be assembled into CSDs or rays, depending on the application. The output, stored as NumPy arrays, can be directly analyzed and plotted, or optionally augmented with noise to emulate experimental data.
@@ 2.1 Nanowire model
QDFlow employs a 1D physics model in which charges are assumed to be confined to a linear nanowire that lies along the $x$-axis. The ends of the nanowire are connected to electron reservoirs, and a bias voltage can be applied between them. Electrostatic gates are positioned at a height $h$ below the xyx\!y-plane, and are modeled as infinite cylindrical conductors with central axis parallel to the $y$-axis, as shown in Fig. 2(a). The arrangement makes our nanowire model a hybrid between a true nanowire device and other QD device architectures. Gates biased to low potential act as plunger gates, while those biased to high potential act as barrier gates (for positive charge carriers, with the convention reversed for negative carriers).
@@
where $V_{h}$ is the potential at a reference distance $r=h$ (chosen as the separation between the gate and the nanowire), is the screening length, and $\mathcal{K}_{0}(z)$ is the modified Bessel function of the second kind. Specifically, we note that $V_{h}$ is not the voltage of the gate itself; rather, it is the voltage that the nanowire feels due to the gate (in the absence of other gates). This is essentially the voltage of the gate multiplied by the lever arm of the gate.
@@
To make the problem tractable, we adopt the simplifying assumption that the induced charges on each gate are rotationally symmetric about the axis of the gate. Under this approximation, the induced charges act as a line charge that lies along the central axis of the gate. Because the gate is rotationally symmetric, the potential $V_{\text{gate}}(r)$ also acts as a single line charge at the center of the gate, and thus the induced charges effectively rescale $V_{\text{gate}}(r)$ by a constant factor. Let $V^{\prime}_{i}$ be the rescaled value of $V_{h}$ for gate $i$ after including the effects of the induced charges on gate $i$, and let $V_{i}$ be the value of $V_{h}$ necessary for $V_{\text{gate}}(r)$ to give the actual potential of gate $i$. Then by the superposition principle, $V_{i}$ can be determined by adding the potential contributions from each of the gates:
$V_{i}=\sum\nolimits_{j}A_{ij}V^{\prime}_{j}$ (2)
@@
$A_{ij}=\begin{cases}1&\text{if }i=j\\ V_{\text{gate}}(x_{j}-x_{i})/V_{\text{gate}}({}_{j})&\text{otherwise,}\end{cases}$ (3)
@@ 2.2 Thomas-Fermi solver
${}^{\prime}(x)=qV(x)+\int_{\mathds{R}}K(x,x^{\prime})n(x^{\prime})dx^{\prime}$ (5)
The physical parameters , $g_{0}$, and in Eq. (4) indicate the Fermi level, the density of states in the conduction band (which is constant in 2D), and the inverse temperature, respectively, and $\operatorname{sp}(z)=\ln(1+e^{z})$ is the softplus function. Parameter $q$ in Eq. (5) controls the sign of the charge carriers, with 1-1 indicating electrons and +1+1 indicating holes, while $K(x,x^{\prime})$ gives the strength of the Coulomb interaction between points $x$ and $x^{\prime}$, and is defined as follows:
$K(x,x^{\prime})=\frac{K_{0}}{\sqrt{(x-x^{\prime})^{2}+{}^{2}}}$ (6)
@@
$\mathbf{K}f(x)=\int_{\mathds{R}}K(x,x^{\prime})f(x^{\prime})dx^{\prime}$ (7)
@@ 2.3 Capacitance model
After calculating $n(x)$, QDFlow employs a capacitance model to determine the stable charge configuration and other properties. Similar techniques have been implemented in other QD simulations [16, 17, 18]. In most of those approaches, the capacitance matrix is assumed to be constant, i.e., the interdot capacitances remain fixed as the gate voltages are swept. The simulation introduced in Ref. 18 allows for variable capacitances by applying a correction to the capacitance matrix based on the particle number. In contrast, QDFlow derives the capacitance matrix directly from the charge density $n(x)$, which depends explicitly on the gate voltages. This feature enables charge-transition slopes and spacings to vary across a single CSD. Moreover, constructing the capacitance model from $n(x)$ naturally captures transitions between a double dot and a merged single dot as the interdot barrier is lowered.
The first step in creating the capacitance model is determining the regions of the nanowire where significant amounts of charge are induced. This is achieved by applying a threshold to $n(x)$, configurable through the NumericsParameters dataclass, and identifying continuous intervals of points that lie above the threshold. This will result in a set of intervals of the form $a_{i},b_{i}]$, which we call “charge islands.” The thresholding is also responsible for determining whether or not adjacent QDs should be handled as individual dots. Specifically, if $n(x)$ exceeds the threshold throughout the region between the two QDs, they are merged and treated as a single dot. Otherwise, they are considered to be two separate QDs with a potential barrier between them.
@@
$E=\sum_{i,j}E_{ij}(Q_{i}-Z_{i})(Q_{j}-Z_{j})$ (13)
$Z_{i}=\int_{a_{i}}^{b_{i}}n(x)dx$ (14)
$E_{ij}=\frac{1}{Z_{i}Z_{j}}\left[c_{k}{}_{ij}\int_{a_{i}}^{b_{i}}n(x)^{2}dx+\frac{1}{2}\int_{a_{i}}^{b_{i}}\int_{a_{j}}^{b_{j}}K(x,x^{\prime})n(x)n(x^{\prime})dxdx^{\prime}\right]$ (15)
@@
After calculating the energy matrix, the next step is to determine the charge configuration $\vec{Q}$ that minimizes the total energy $E$, subject to the constraint that all $Q_{i}$ must be nonnegative integers. This is an instance of an integer optimization problem, which in general is NP-complete. However, for a moderate number of gates, a brute-force search is sufficient to find the minimum. In particular, we first locate the minimum in the continuous space, which occurs at $\vec{Z}$, and then evaluate $E(\vec{Q})$ over all $\vec{Q}$ such that for each integer $Q_{i}$, $|Q_{i}-Z_{i}|<1$. Once a stable charge configuration is identified, the potential at each of the sensors is calculated under the assumption that each island $i$ hosts a line of charge with total charge $q\,Q_{i}$ and charge density proportional to $n(x)$. The Coulomb potential at each of the sensors arising from these charge islands is calculated, and the result is normalized by dividing by the potential of a single point charge located at a point on the nanowire closest to the sensor in question. This means that a single transition should have a height of no more than 1 after normalization.
Finally, QDFlow allows the current across the nanowire to be found. For this calculation, the left and right sides of the nanowire are assumed to be connected to electron baths with potentials $V_{L}$ and $V_{R}$, respectively. The dynamics of the charges are modeled using a semi-classical approach, treating them as particles that travel at the Fermi velocity. Each time they collide with a barrier, the particles have a chance to either tunnel through it or be reflected back. The tunneling probability across each barrier is determined by the transmission coefficient, which we calculate using the WKB approximation. This allows the tunneling rates between islands and the tunneling rates to and from the external charge baths to be obtained. These tunnel rates are then used to define a Markov graph which encodes the dynamics of the transitions between charge states. The current through the nanowire is obtained by evaluating the net rate at which charges enter and leave the charge baths at the steady state of this Markov graph.

### v3

# QDFlow: A Python package for physics simulations of quantum dot devices

@@ Abstract
Recent advances in machine learning (ML) have accelerated progress in calibrating and operating quantum dot (QD) devices. However, most ML approaches rely on access to large, representative datasets designed to capture the full spectrum of data quality encountered in practice, with both high- and low-quality data for training, benchmarking, and validation, with labels capturing key features of the device state. Collating such datasets experimentally is challenging due to limited data availability, slow measurement bandwidths, and the labor-intensive nature of labeling. QDFlow is an open-source physics simulator for multi-QD arrays that generates realistic synthetic data with ground-truth labels. QDFlow combines a self-consistent Thomas-Fermi solver, a dynamic capacitance model, and flexible noise modules to simulate charge stability diagrams and ray-based data that closely resemble experimental results. With an extensive set of parameters that can be varied and customizable noise models, QDFlow supports the creation of large, diverse datasets for ML development, benchmarking, and quantum device research.
@@ 1 Introduction
Among the various quantum computing platforms, quantum dots (QDs) stand out for their scalability potential, compact size, and long coherence times [1]. Operating QD devices, however, remains a formidable challenge, with complexity growing rapidly—often exponentially—as the number of qubits increases. Recent advances in integrating machine learning (ML) with quantum device operation have begun to mitigate these difficulties, offering promising automated control and calibration strategies. For example, ML algorithms have been developed for the fabrication [2, 3], characterization [4, 5], tuning [6, 7, 8, 9, 10, 11, 12, 13], and gate virtualization [14, 15] of QD devices.
Developing robust ML models requires access to large, diverse datasets that are representative of the multi-dimensional parameter space typical of QD devices. Crucially, for supervised ML applications, these datasets must also include metadata that identifies key features, such as the global state (i.e., the number of QDs formed), the charge configuration, and the type of transition lines present. Unfortunately, large volumes of high-quality experimental data can be challenging to obtain as companies and research groups often keep such data proprietary [16]. Limited measurement bandwidth in real-world experiments also constrains the efficient exploration of the entire high-dimensional parameter space in a reasonable time. Generating accurate feature labels for publicly available data is a labor-intensive and time-consuming task that can yield subjective and potentially erroneous labels. Physics-based simulations offer a practical solution: they enable the generation of arbitrarily large datasets while providing direct access to the ground-truth charge states, thereby simplifying the labeling process needed for ML training.
@@
There are several open-source QD device simulators that rely on the constant capacitance model, treating the array of QDs and their associated electrostatic gates as nodes in a network of fixed capacitors [17, 18, 19]. Additionally, Ref. [19] allows the capacitances to vary with respect to the number of charges $n$ by introducing an $n$-dependent correction to the capacitance matrix. In contrast, in QDFlow the capacitance parameters are physics-informed, obtained directly from the self-consistent Thomas-Fermi solution rather than imposed heuristically. All key physical observables—such as current, charge states, and sensor readouts—are derived from a capacitance model constructed based on the computed charge density $n(x)$. This ensures that the capacitances evolve dynamically with gate voltages, yielding a more realistic description of device behavior. Furthermore, QDFlow allows modeling regions with low barriers between dots, leading the dots on either side to coalesce into a single, centralized dot. Finally, QDFlow incorporates a flexible noise module, enabling the addition of experimentally relevant effects such as thermal broadening, charge offset drift, and voltage fluctuations. These features make the simulated data qualitatively comparable to experimental measurements while maintaining full access to the ground truth labels required for ML applications. Building on QFlow—a legacy implementation of the QD simulator that applied the Thomas-Fermi approximation to model charge densities and stability diagrams [20, 21]QDFlow extends these methods into a flexible, open-source framework that integrates physics-informed capacitance modeling with realistic noise processes tailored for ML applications.
@@
QDFlow has already demonstrated its utility in advancing ML-driven QD research, confirming that ML models trained exclusively on simulated data can be successfully deployed in experimental settings, on samples fabricated in an academic cleanroom as well as on an industrial 300-$\mathrm{mm}$ process line, and on 1D and 2D QD arrays. The legacy version of the simulator was used to generate the QFlow-lite dataset [22, 21], which enabled the training of several ML models for global state recognition. The utility of these ML models was demonstrated both offline, by navigating the voltage space within pre-measured experimental datasets [20, 13], and in closed-loop experiments [9, 12, 23]. The dataset also supported the development of a novel classification framework for simple high-dimensional geometrical structures, known as the ray-based classification (RBC) framework [24]. The expanded dataset, QFlow 2.0: Quantum dot data for machine learning [22], generated using the Thomas-Fermi solver with integrated realistic noise processes, further advanced ML-based approaches to QD tuning. In particular, models trained with data from the QFlow 2.0 dataset have been successfully applied to tasks such as data quality assessment[25], physics-informed RBC and ray-based navigation in 1D QD arrays [13], and the development of a full virtualization stack for 1D and 2D QD arrays in Ge/SiGe [14] and in Si/SiGe. These successes confirm the compatibility of QFlow 2.0-generated with real-world experiments. They also highlight the value of physics-informed synthetic datasets for accelerating the development of automated control tools for QD systems.
@@
QDFlow is available for download from the Python Package Index, with the source code released under the GNU General Public License in the QDFlow GitHub repository [26]. Comprehensive API documentation is provided via docstrings embedded in the source code and as HTML pages hosted on Read the Docs. The library includes type hints for all classes and functions to support clarity, maintainability, and extensibility. Unit and benchmark tests are also distributed with the QDFlow repository to facilitate validation and performance evaluation.
@@ 2 Physics simulation
QDFlow has three main modules: simulation, noise, and generate, and one utility module, distribution, as depicted in Fig. 1. The simulation and noise modules are part of the physics package. The core physics-based engine of the simulator is contained in the simulation module. It uses a Thomas-Fermi solver to find the stable charge configuration and sensor output of a particular QD device defined by a set of physical parameters. The QDFlow Thomas-Fermi solver was first introduced in Ref. [21], but has since been refined and extended within QDFlow to improve flexibility, physical relevance, and integration with downstream ML workflows. The PhysicsParameters dataclass, which is used to initialize the simulation, specifies over twenty parameters governing the properties of the QD device. These parameters include both material characteristics and device-specific features such as gate geometry and positioning. Importantly, the gate voltages—experimentally relevant control knobs—are explicitly included among the simulation inputs. By sweeping these voltages, QDFlow produces the final outputs: 2D CSDs and 1D rays, directly mirroring the tuning procedures used in real QD experiments.
@@
To generate data with QDFlow, the user first chooses whether to run the default configuration or adjust the distributions and ranges used to randomize the physics parameters. Next, they create one or more sets of randomized device parameters, and for each device, generate a CSD using the functions in the generate module. Once the physics parameters have been specified, an instance of the ThomasFermi class is instantiated. This class serves two main purposes: first, it solves for the charge density function $n(x)$; and second, it uses $n(x)$ to construct a capacitance model and compute physical quantities such as the device’s charge state and sensor response. QDFlow then runs the physics simulation for every pixel in each diagram and compiles the results into a CSDOutput dataclass, which is returned to the user. By repeating this process over a range of gate voltages, QDFlow generates data that can be assembled into CSDs or rays, depending on the application. The output, stored as NumPy arrays, can be analyzed and plotted directly and, optionally, augmented with noise to emulate experimental data.
@@ 2.1 Nanowire model
QDFlow employs a 1D physics model in which charges are assumed to be confined to a linear nanowire that lies along the $x$-axis. The ends of the nanowire are connected to electron reservoirs, and a bias voltage can be applied between them. Electrostatic gates are positioned at a height $h$ below the xyx\!y-plane, and are modeled as infinite cylindrical conductors with central axis parallel to the $y$-axis, as shown in Fig. 2(a). The arrangement makes our nanowire model a hybrid between a true nanowire device and other QD device architectures. Gates biased to low potential act as plunger gates, while those biased to high potential act as barrier gates (for positive charge carriers; the convention is reversed for negative carriers).
@@
where $V_{h}$ is the potential at a reference distance $r=h$ (chosen as the separation between the gate and the nanowire), is the screening length, and $\mathcal{K}_{0}(z)$ is the modified Bessel function of the second kind. Specifically, we note that $V_{h}$ is not the gate voltage itself; rather, it is the voltage the nanowire experiences due to the gate (in the absence of other gates). This is essentially the gate voltage multiplied by the gate lever arm.
@@
To make the problem tractable, we adopt the simplifying assumption that the induced charges on each gate are rotationally symmetric about the axis of the gate. Under this approximation, the induced charges act as a line charge along the gate’s central axis. Because the gate is rotationally symmetric, the potential $V_{\text{gate}}(r)$ also acts as a single line charge at the center of the gate, and thus the induced charges effectively rescale $V_{\text{gate}}(r)$ by a constant factor. Let $V^{\prime}_{i}$ be the rescaled value of $V_{h}$ for gate $i$ after including the effects of the induced charges on gate $i$, and let $V_{i}$ be the value of $V_{h}$ necessary for $V_{\text{gate}}(r)$ to give the actual potential of gate $i$. Then, by the superposition principle, $V_{i}$ can be determined by adding the potential contributions from each of the gates:
$V_{i}=\sum\nolimits_{j}A_{ij}V^{\prime}_{j},$ (2)
@@
$A_{ij}=\begin{cases}1,&\text{if }i=j,\\ V_{\text{gate}}(x_{j}-x_{i})/V_{\text{gate}}({}_{j}),&\text{otherwise,}\end{cases}$ (3)
@@ 2.2 Thomas-Fermi solver
${}^{\prime}(x)=qV(x)+\int_{\mathds{R}}K(x,x^{\prime})n(x^{\prime})dx^{\prime}.$ (5)
The physical parameters , $g_{0}$, and in Eq. (4) indicate the Fermi level, the density of states in the conduction band (which is constant in 2D), and the inverse temperature, respectively, and $\operatorname{sp}(z)=\ln(1+\exp(z))$ is the softplus function. Parameter $q$ in Eq. (5) controls the sign of the charge carriers, with 1-1 indicating electrons and +1+1 indicating holes, while $K(x,x^{\prime})$ gives the strength of the Coulomb interaction between points $x$ and $x^{\prime}$, and is defined as follows:
$K(x,x^{\prime})=\frac{K_{0}}{\sqrt{(x-x^{\prime})^{2}+{}^{2}}},$ (6)
@@
$\mathbf{K}f(x)=\int_{\mathds{R}}K(x,x^{\prime})f(x^{\prime})dx^{\prime}.$ (7)
@@ 2.3 Capacitance model
After calculating $n(x)$, QDFlow employs a capacitance model to determine the stable charge configuration and other properties. Similar techniques have been implemented in other QD simulations [17, 18, 19]. In most of those approaches, the capacitance matrix is assumed to be constant, i.e., the interdot capacitances remain fixed as the gate voltages are swept. The simulation introduced in Ref. [19] allows for variable capacitances by applying a correction to the capacitance matrix based on the particle number. In contrast, QDFlow derives the capacitance matrix directly from the charge density $n(x)$, which depends explicitly on the gate voltages. This feature enables charge-transition slopes and spacings to vary across a single CSD. Moreover, constructing the capacitance model from $n(x)$ naturally captures transitions between a double dot and a merged single dot as the interdot barrier is lowered.
The first step in creating the capacitance model is to determine the regions of the nanowire where significant charge is induced. This is achieved by applying a threshold to $n(x)$, configurable through the NumericsParameters dataclass, and identifying continuous intervals of points that lie above the threshold. This will result in a set of intervals of the form $a_{i},b_{i}]$, which we call “charge islands.” Thresholding also determines whether adjacent QDs should be treated as individual dots. Specifically, if $n(x)$ exceeds the threshold throughout the region between the two QDs, they are merged and treated as a single dot. Otherwise, they are considered to be two separate QDs with a potential barrier between them.
@@
$E=\sum_{i,j}E_{ij}(Q_{i}-Z_{i})(Q_{j}-Z_{j}),$ (13)
$Z_{i}=\int_{a_{i}}^{b_{i}}n(x)dx,$ (14)
$E_{ij}=\frac{1}{Z_{i}Z_{j}}\left[c_{k}{}_{ij}\int_{a_{i}}^{b_{i}}n(x)^{2}dx+\frac{1}{2}\int_{a_{i}}^{b_{i}}\int_{a_{j}}^{b_{j}}K(x,x^{\prime})n(x)n(x^{\prime})dxdx^{\prime}\right],$ (15)
@@
After calculating the energy matrix, the next step is to determine the charge configuration $\vec{Q}$ that minimizes the total energy $E$, subject to the constraint that all $Q_{i}$ must be nonnegative integers. This is an instance of an integer optimization problem, which in general is NP-complete. However, for a moderate number of gates, a brute-force search is sufficient to find the minimum. In particular, we first locate the minimum in the continuous space, which occurs at $\vec{Z}$, and then evaluate $E(\vec{Q})$ over all $\vec{Q}$ such that for each integer $Q_{i}$, $|Q_{i}-Z_{i}|<1$. Once a stable charge configuration is identified, the potential at each of the sensors is calculated under the assumption that each island $i$ hosts a line of charge with total charge $q\,Q_{i}$ and charge density proportional to $n(x)$. The Coulomb potential at each sensor arising from these charge islands is calculated, and the result is normalized by dividing by the potential of a single point charge at the point on the nanowire closest to the sensor in question. This means that a single transition should have a height of no more than 1 after normalization.
Finally, QDFlow allows the current across the nanowire to be found. For this calculation, the left and right sides of the nanowire are assumed to be connected to electron baths with potentials $V_{L}$ and $V_{R}$, respectively. The dynamics of the charges are modeled using a semiclassical approach, treating them as particles that travel at the Fermi velocity. Each time they collide with a barrier, the particles have a chance to either tunnel through it or be reflected back. The tunneling probability across each barrier is determined by the transmission coefficient, which we calculate using the WKB approximation. This allows the tunneling rates between islands and to and from the external charge baths to be obtained. These tunnel rates are then used to define a Markov graph which encodes the dynamics of the transitions between charge states. The current through the nanowire is obtained by evaluating the net rate at which charges enter and leave the charge baths at the steady state of this Markov graph.